How many moles of H3O+ or OH− must you add to a liter of strong acid solution to adjust its pH from 3.46 to 3.96? Assume a negligible volume change.
Answer #1

The formula for pH is as follows:

pH = -log[H3O+]

[H3O+] = 10^-pH

Initial pH = 3.46

Then [H3O+] = 10^-3.46 = 3.467×10^-4 M

Final pH = 3.96

Then [H3O+] = 10^-3.96 = 1.096×10^-4 M

The change in concentration of H3O+ is

3.467×10^-4 M - 1.096×10^-4 M = 2.371 ×10^-4 M

This change is equal to concentration of base OH-.

Hence the number of moles of OH- added is 2.37×10^-4 or 0.000237 moles